OpenStudy (anonymous):

For every _____ degrees equal 1 radian. I forgot the rule:/

OpenStudy (acxbox22):

http://www.mathsisfun.com/geometry/radians.html

OpenStudy (agreene):

its in the neighbourhood of 57.295

OpenStudy (anonymous):

Thanks @acxbox22 and @agreene

OpenStudy (acxbox22):

np

OpenStudy (anonymous):

So if I have 82 degrees how would i find the number of radians?

OpenStudy (acxbox22):

divide it by 57.295

OpenStudy (agreene):

btw: \[1 \space \text{rad}\approx \frac{180°}{\pi}\]

OpenStudy (anonymous):

So then I get 1.43118946. Thanks again!

OpenStudy (agreene):

\[\text{angle in radians}=\text{angle in degrees}*\frac{\pi}{180}\]